The Ramapo High School lacrosse team sold brownies and cookies to earn money for a tournament. The brownies sold for $6.00 per box and the cookies sold for $8.00 a box. The total number of baked goods was 580. and the total amount of money collected was $4,340. How many boxes of each kind of baked goods were sold?
Answer by ewatrrr(24785)


Hi,
Let x represent the number of $8 boxes. then (580-x) would be the number of $6 boxes
8x + 6(580-x) = $4340
2x = 4340 - 3480
2x = 860
x = 430, the number of $8 boxes. 150 the number of $6 boxes (580-430)
checking our answer
$8*430 + $6*150 = $4340